Given the supply function:
Q = 150 + 5P + 0.1√(2)
Find price elasticity of supply when P = 10
1
Expert's answer
2018-04-13T10:17:11-0400
Given, supply function Q = 150 + 5P + 0.1 √(2) ……………………..(1)
At P= 10 equation (1) becomes, Q= 150 + (5×10) +0.1 √(2) = 200.14
Now differentiate equation (1) with respect to P , we get dQ/dP=5
We know price elasticity of supply (e) = dQ/dP ×P/Q = 5×P/Q
Now at P = 10, price elasticity of supply (e)= 5×10/200.14 = 0.2498 (approx.)
(put the value of P and Q at P = 10)


Answer: Price elasticity of supply is 0.2498.